OpenStudy (arthur_ser):

High school math. Prove that:

OpenStudy (arthur_ser):

\[2\int\limits_{a}^{b}f(x) f \prime(x)dx= (f(b))^{2}-(f(a))^{2}\]

OpenStudy (arthur_ser):

Any help would be appreciated

OpenStudy (anonymous):

my first impulse would be the u - sub \(u=f(x)\)

OpenStudy (anonymous):

should work exactly you good from there?

OpenStudy (arthur_ser):

So the left hand-side becomes: \[2\int\limits_{f(a)}^{f(b)}u \times u \prime du \] right?